OpenStudy (shesolitt):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form? I WILL MEDAL AND FAN ! y = 3x + 2 y = 3x + 8 y = –3x – 2 y = –3x + 8

OpenStudy (pawanyadav):

m=(14-5)/(-2-1)=-3 And y=mx +c 5=-3+c c=8 So D is correct

OpenStudy (shesolitt):

TY !!!